OpenStudy (anonymous):

7( x - 1 ) - x = y. x = 5?

OpenStudy (anonymous):

if x = 5 then 7( 5 - 1 ) - 5, do whats in the parenthesis first

OpenStudy (anonymous):

7(x-1) - x = y Given x=5, so just put the value in the given equation. i.e. \[y = 7(5-1) -5\]Using the BODMAS rule, i.e. first work with parentheses. \[y = 7\times 4 -5\]Finally,\[y = 28-5 = 23\]